you put 5000$ in the bank the beginning of the year. From there, you earn 10% interest each year. how much money do you make after 50 years?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(12270)

The formula for amount =p(1+(r/100))^n
p=50,000
r=10
n=50 years
Amount = 50000(1+10/100)^50
= 50000(11/10)^50
=$5,869,542.64 will be the amount after 50 years

Answer by richwmiller(9143)

that's 5000 not 50,000 for p
If his formula is right that yields
$586,954.24